OpenStudy (anonymous):

y = (e^x + 4^3x)^6 find the derivative

OpenStudy (anonymous):

is it 6(e^x+4^3x)^5∗(2^ex+ln(4)(3)+4^3x)

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

the derivative of \[4^{3x}\] is \[3\times \ln(4)\times 4^{3x}\] by the chain rule

OpenStudy (anonymous):

you have plus. it should be times